Please help me to find the value of x for which fraction x^2-9/2x^2-7x+3 is undefine
Answer by stanbon(75887)

Please help me to find the value of x for which fraction
x^2-9/2x^2-7x+3 is undefined.
------
Undefined when the denominator = 0
--------
Solve: 2x^2 - 7x + 3 = 0
----
x = [7 +- sqrt(49-4*2*3)]/4
------
x = [7 +- sqrt(25)]/4
---
x = (7+5)/4 or x = (7-5)/4
----
x = 3 or x = 1/2
------------------------
Cheers,
Stan H.
========================